OpenStudy (anonymous):

Use the matrix tool to solve the system of equations. Answer as ordered pair. 3x-8y=-1 -2x+6y=1

OpenStudy (anonymous):

I am unfamiliar with the Cramer rule as well as the matrix. Kinda new for me. @uri

OpenStudy (uri):

So how will we solve it..if we don't know..wait elimination?substitution?

OpenStudy (anonymous):

I believe its elimination @uri

OpenStudy (uri):

3x-8y=-1(multiply by 3) -2x+6y=1(multiply by 4) 6x-24y=-3 -8x+24y=4 ------------ -2x=1 x=1/-2

OpenStudy (uri):

To find y use the original 2nd equation and put the value of x -2x+6y=1 -2(1/-2)+6y=1 1+6y=1 subract 1 both sides 1+6y=1 -1 -1 ====== 6y ans.

OpenStudy (anonymous):

So as an ordered pair it would be (0.5,6)?

OpenStudy (anonymous):

@uri

OpenStudy (uri):

Yes -0.5,6

OpenStudy (anonymous):

Thank you @uri